OpenStudy (anonymous):

Would someone help me with this problem? How many 160 resistors must be connected in parallel to draw a current of 6A from a 120V source?

OpenStudy (roadjester):

R=V/I so R=120V/6A which means that you want an equivalent resistance of 20 ohms

OpenStudy (roadjester):

1/R_eq=(1/R1)+(1/R2)+(1/R3)+...+(1/Rn) where n is the number of resistors. But R1=R2=R3=...Rn=160 ohms

OpenStudy (roadjester):

 and "n" is the number of resistors